### v1

@@ Abstract
The entertainment-driven dynamics of social media platforms encourage users to engage with like-minded individuals and consume content aligned with their beliefs. These dynamics may amplify polarization by reinforcing shared perspectives and reducing exposure to diverse viewpoints. Simultaneously, users migrate from one platform to another, either forced by moderation policies, such as de-platforming, or spontaneously seeking environments more aligned with their preferences. These migrations foster the specialization and differentiation of the social media ecosystem, with platforms increasingly organized around specific user communities and shared content preferences. This shift marks an evolution from echo chambers enclosed within platforms to “echo platforms”, i.e., entire platforms functioning as ideologically homogeneous niches. This study introduces an operational framework to systematically analyze these dynamics, by examining three key dimensions: platform centrality (central vs. peripheral), news consumption (reliable vs questionable), and user base composition (uniform vs diverse). To this aim, we leverage a dataset of 126M URLs posted by nearly 6M users on nine social media platforms, namely Facebook, Reddit, Twitter (now X), YouTube, BitChute, Gab, Parler, Scored, and Voat. We find a clear separation between mainstream and alt-tech platforms, with the second category being characterized by a peripheral role in the social media ecosystem, a greater prevalence of unreliable content, and a heightened ideological uniformity. These findings outline the main dimensions defining the fragmentation and polarization of the social media ecosystem.
Introduction
@@
Social media are now a major source of news and opinions for many, reshaping how people access information and engage in public discussions. This shift comes as trust in traditional media continues to decline [1, 2]. This transformation is driven by platforms prioritizing entertainment and user engagement over informational accuracy [3, 4, 5, 6]. Such dynamics, combined with individual preferences, encourage users to join like-minded peers and consume content aligning with their pre-existing beliefs, reinforcing shared perspectives and potentially limiting exposure to diverse viewpoints [7, 5, 8]. These behaviors have been closely linked to heightened exposure to hate speech [9, 10, 11, 12], political polarization [13, 14, 15], and the formation of echo chambers [5, 16]. Such clusters amplify ideological homogeneity, further reduce exposure to opposing viewpoints, and shape opinions through repeated interactions [17, 18]. Over time, this clustering may drive groups toward more extreme positions [19, 20], influencing user behavior across platforms [21, 5, 22, 23]. Beyond digital interactions, this polarization affects offline social cohesion and public discourse, highlighting its broader societal consequences.
To address these challenges, platforms have increasingly implemented moderation policies aimed at managing tight-knit, “problematic” communities [24]. However, moderation policies, such as de-platforming, often push users toward minimally regulated platforms [25, 16, 26, 27, 28]. At the same time, users may spontaneously leave platforms where they find limited alignment with their preferences, seeking alternatives. For example, many users migrated from Twitter to Mastodon [29] or BlueSky [30, 31] following changes in platform management.
When users leave mainstream platforms, they often migrate to less popular digital spaces, commonly referred to as “alt-tech” or “fringe” platforms [32, 33, 26]. Examples include Gab, Parler, BitChute, and Rumble, which have emerged as counterparts to mainstream services like Twitter, Reddit, YouTube, and Facebook. These platforms frequently market themselves as champions of free speech, appealing to users who perceive mainstream platforms as ideologically biased or overly regulated. Such migrations, whether driven by dissatisfaction with moderation policies or forced exclusion through bans, contribute to the fragmentation of the digital public sphere [34, 35].
@@
Both spontaneous and forced migrations foster differentiation and specialization within the social media ecosystem, with platforms increasingly organized around specific user communities and shared content preferences [36, 37, 38, 39]. These shifts represent a departure from the internet’s original vision as a globally interconnected network where information flows freely across borders [40, 41]. Instead, entire platforms are now defined by ideologically homogeneous communities and aligned content, creating niches where opposing perspectives are rarely encountered [37]. This scenario—where an entire platform is characterized by a like-minded user base with limited exposure to diverse opinions—lays the foundation for the concept of an “echo platform”, an extension of the echo chamber phenomenon to the level of entire digital ecosystems.
@@
In this paper, we introduce an operational framework to characterize the fragmentation of the social media ecosystem, providing a methodological basis for identifying the role of different platforms. Using a dataset comprising over 126 million URLs collected from nine platforms—four mainstream (Facebook, Twitter, Reddit, YouTube) and five alt-tech or fringe platforms (BitChute, Gab, Parler, Scored, Voat)—and nearly six million unique users, we define three axes to characterize platform roles: (i) centrality (central vs. peripheral), (ii) news consumption (reliable vs. questionable content), and (iii) user base composition (uniform vs. diverse).
@@
Our analysis reveals distinct patterns of fragmentation and information consumption. Interactions between mainstream and alt-tech platforms are limited, reinforcing the fragmentation of the social media ecosystem. Alt-tech platforms exhibit significantly higher levels of ideological homogeneity, with user communities sharing content almost exclusively aligned with their dominant narratives. Interestingly, user behavior on Reddit more closely resembles alt-tech platforms than mainstream ones. Furthermore, alt-tech platforms disproportionately amplify questionable content while showing a notable absence of reliable news sources compared to their mainstream counterparts.
@@ Results
The systematic characterization of these three dimensions provides a robust framework for understanding the role of different platforms in shaping the digital public sphere. Moreover, it underscores areas for targeted policy and research interventions to mitigate polarization and misinformation.
The characterization is based on a comprehensive dataset of URLs collected from posts (or video descriptions for YouTube and BitChute) across nine social media platforms. As shown in Table 1, this dataset includes four “mainstream” platforms (Facebook, Twitter, Reddit, and YouTube) and five “alt-tech” platforms (BitChute, Gab, Parler, Scored, and Voat). These URLs, linking to internal and external domains, serve as a proxy for a macro-level analysis of platform behavior (see the Materials and Methods section for further details).
@@
Facebook 300k 10.7M 74k 0.09 0.15 0.16
Reddit 59k 320k 9k 0.14 0.03 0.04
Twitter 5M 103M 114k 0.16 0.16 0.16
YouTube 12k 512k 12k 0.11 0.14 0.19
BitChute 17k 1.9M 45k 0.14 0.17 0.21
Gab 15k 453k 7k 0.08 0.65 0.06
Parler 300k 8.2M 70k 0.11 0.51 0.07
Scored 45k 883k 12k 0.06 0.77 0.04
Voat 27k 215k 10k 0.12 0.36 0.11
@@
To analyze the roles of different platforms within the information ecosystem, we model it as a weighted, directed graph. In this network, nodes represent platforms, and the weight of a directed edge between two nodes corresponds to the number of URL links pointing from one platform to another. In other words, an edge from platform $i$ to platform $j$ exists if an account of platform $i$ posted a URL linking directly to platform $j$.
As the nodes of the network correspond to the platforms we take into account, a full representation of the weighted adjacency matrix (see SI) can provide insights into how platforms redirect users and share content within the ecosystem. However, to account for variations in the size of the datasets collected from each platform, we display the observed edge weights rescaled by those generated using a null model that preserves the total number of outgoing and incoming URL links for each platform (see section Materials and Methods for details).
@@
Fig. 1 displays the ratio of observed edge weights to those expected under the null model. Ratios greater than one indicate that a platform links to another more frequently than expected under random distribution, while ratios less than one indicate less frequent linking. Fig. 1 reveals distinct patterns of inter-platform linking: Twitter consistently receives more URLs than expected from all other platforms, positioning it as a central node in the information ecosystem. Surprisingly, Facebook receives fewer links than expected from other mainstream platforms, except YouTube, likely due to the prevalence of self-promotional links in video descriptions. Alt-tech platforms generally receive fewer links than expected by mainstream ones, and are instead more tightly connected among themselves. Nevertheless, the strongest link weight with respect to the null model is observed from Gab to Reddit, suggesting a potential content-sharing dynamic or migration pathway between these two platforms. BitChute is linked much more than expected by all fringe platforms and slightly less by mainstream ones, while the opposite is true for YouTube. Finally, Scored receives no incoming links from six of the nine platforms, likely due to its limited popularity and low exposure, underscoring its peripheral role in the broader ecosystem.
To better quantify the relative importance of each platform within the ecosystem, we computed their PageRank centrality from the rescaled adjacency matrix shown in Fig. 1. The results presented in Table 1 reveal that Twitter, Reddit, and BitChute exhibit the highest scores, reflecting their prominent roles in the network. Voat, Parler, and YouTube display lower centrality scores but still maintain relatively strong connection patterns. Facebook, Gab, and Scored rank as the least central platforms in the ecosystem.
@@
Next, we investigate whether the type of news and content shared differs significantly across the nine platforms by analyzing the news domains’ political bias. To this aim, we use Media Bias/Fact Check (commonly referred to as MBFC, which can be accessed by https://mediabiasfactcheck.com), a widely referenced news rating agency that provides political bias and reliability labels for news outlets and other information producers [42, 5, 43]. We extract the domains from the URLs shared across the nine platforms and match them with the political bias and reliability information obtained from MBFC. Domains not classified by MBFC are labeled as “unreported”, though this designation does not necessarily imply an absence of political bias. Domains categorized as “extreme-left” are excluded from the analysis due to their negligible presence in the dataset (see Materials and Methods for details).
Fig. 2 illustrates users’ news consumption across platforms, showing the proportion of URLs directing to domains with differing political leanings, as categorized by MBFC. A clear pattern emerges: mainstream platforms such as Facebook, Reddit, and Twitter tend to skew toward left-leaning content, while alt-tech platforms—including Gab, Parler, Voat, and Scored—predominantly link to right-leaning content. Voat stands out among alt-tech platforms by directing a notable portion of its traffic to left- or left-center-leaning domains. Interestingly, 79.7% of these links were shared across eleven of the 2,423 distinct subverses analyzed, many of which (e.g., v/AnonAll, v/GreatAwakening) align with alt-right or conspiracy groups (a detailed list of subverses is available in the SI). This pattern suggests that these links may have been shared in an ironic or derogatory manner. YouTube and BitChute heavily link to “unreported” domains, reflecting their role as platforms that host a broad range of less conventional or unvetted content.
@@
To further illustrate these patterns, Table 2 reports each platform’s ten most frequently linked domains. While Facebook and Twitter often feature sources having a left-center bias, such as CNN, The New York Times, or The Washington Post, Reddit displays a more homogenous profile as no right-biased sources appear among the most linked domains. In contrast, platforms such as Gab, Parler, Scored, and Voat direct a substantial proportion of their traffic to sources that are both right-biased and have low reliability. Scored, for instance, channels 55% of its external traffic to “Patriots.Win”, an online forum created following the ban of the “r/The_Donald” subreddit in June 2015 [33]. This forum is known for using unreliable sources and a lack of transparency and moderation [44, 45]. Although both YouTube and BitChute frequently link to “unreported” domains, BitChute’s links often direct users to alternative news websites, including extremist domains like “Glavasic.com”, a platform created by Serbian philosopher Dragan Glavašić to promote his white supremacist manifesto.
@@
Facebook Reddit Twitter YouTube BitChute Domain % links Domain % links Domain % links Domain % links Domain % links CNN 2.35 The Hill 6.00 The New York Times 6.90 Fox News* 8.09 Infowars* 2.12 Fox News* 2.30 CNN 5.82 The Washington Post 6.56 MSNBC 1.97 Renegade Tribune* 1.82 The Hill 1.79 The Washington Post 4.73 CNN 5.72 Tv9Hindi 1.86 Rebel News* 1.36 The Washington Post 1.72 The New York Times 3.95 Fox News* 3.22 The Young Turks 1.77 London Real 1.13 The New York Times 1.65 Politico 2.79 Politico 2.34 CNBC 1.55 Zero Hedge* 0.81 NBC News 1.57 Reuters 2.37 The Hill 2.30 PBS NewsHour 1.47 Brave 0.74 Yahoo News 1.50 Associated Press 2.18 NBC News 2.27 Fox Business 1.47 Banned News* 0.74 Breitbart* 1.46 The Guardian 2.10 Breitbart* 2.23 NBC News 1.10 Glavasic.com* 0.69 Daily Wire 1.22 NBC News 2.07 Raw Story 2.19 France 24 0.91 NewsWars* 0.48 MSN 1.04 Business Insider 1.93 The Gateway Pundit* 1.92 CBS News 0.67 fuckthejews* 0.39 Gab Parler Scored Voat Domain % links Domain % links Domain % links Domain % links The Gateway Pundit* 13.71 The Gateway Pundit* 8.35 Patriots* 57.68 The Gateway Pundit* 5.47 Breitbart* 12.07 Fox News* 7.02 Breitbart* 3.36 Breitbart* 4.62 Left Fox News* 4.09 Breitbart* 5.82 MAGA* 2.62 Zero Hedge* 4.05 Left Center Zero Hedge* 2.47 The Epoch Times* 4.45 The Gateway Pundit* 2.31 DailyMail* 2.34 Center The Epoch Times* 2.43 Western Journal* 1.68 Fox News* 1.64 Fox News* 2.12 Right Center InfoWars* 2.29 New York Post 1.63 Zero Hedge* 0.99 RT News* 1.61 Right GNews* 1.97 Town Hall* 1.18 New York Post 0.92 The Hill 1.50 Extreme Right New York Post 1.70 The Blaze* 1.07 Washington Examiner* 0.72 New York Post 1.44 Unreported The Hill 1.31 Zero Hedge* 1.07 DailyMail* 0.70 CNN 1.17 DailyCaller 1.19 The Federalist* 1.07 DailyCaller 0.56 Washington Examiner 1.03
@@
Next, we analyze the similarity between the news diets across the platforms in the information ecosystem. To this aim, we identify the top 20 most-linked domains for each platform and compute the weighted cosine similarity between the domain vectors for all pairs of platforms. This method quantitatively measures how closely platforms align in their shared content. The results, illustrated in Fig. 3, reveal two distinct cliques characterized by high similarity scores. The first clique consists of the mainstream platforms Facebook, Twitter, and Reddit, all exhibiting cosine similarity scores above $0.7$. The second clique includes the alt-tech platforms Gab, Voat, and Parler, with similarity scores ranging from $0.74$ to $0.88$. Conversely, the similarity between mainstream and alt-tech platforms does not exceed $0.5$, with Facebook and Twitter showing slightly higher similarity to specific fringe platforms. This observation may indicate that the news consumption of certain echo chambers on Facebook and Twitter partially overlaps with those on fringe platforms. In contrast, Reddit appears to be the mainstream platform most dissimilar to alt-tech platforms, exhibiting the lowest similarity in news consumption. This aligns with the tendency of Reddit users to predominantly share left-leaning content, while extreme-right sources are absent and right-center or right-leaning content is relatively uncommon. YouTube shows moderate similarity to other mainstream platforms, particularly Twitter, and to Parler for what concerns the alt-tech platforms, while BitChute and Scored are dissimilar to all other platforms. This difference may be attributed to the extreme nature of the content frequently shared on BitChute and the limited popularity and reach of Scored, which results in unique news diets that diverge significantly from both mainstream and other alt-tech platforms. Refer to the SI for the cosine similarity matrix.
@@
Fig. 3 also highlights the fraction of questionable domains (as defined by MBFC) shared by each platform. Mainstream platforms share a relatively small fraction of unreliable sources, while alt-tech platforms display significantly higher proportions of such content, reinforcing the distinct content dynamics within the information ecosystem. These results are reported in more detail in Table 1, showing the fraction of questionable sources shared by users on each platform. Alt-tech platforms Gab, Parler, and especially Scored exhibit the highest proportions. Voat shows a lower fraction, while BitChute has a value of $0.17$, similar to the one observed in mainstream platforms such as Facebook, Twitter, and YouTube. Reddit stands out for its extremely low fraction of questionable sources shared ($0.03$). These findings underscore the distinction between alt-tech platforms, where questionable sources appear to circulate more freely, and mainstream platforms, where such content constitutes only a small proportion of shared material. The BitChute results, while differing from the general trend observed among alt-tech platforms, align with the platform’s propensity to host a significant volume of “unreported” domains.
@@
Figure 3: Cosine similarity network based on the platforms’ 20 most linked domains. The size of the different nodes is proportional to the volume of links shared in the platform, while the colors of the pies indicate the fraction of questionable or reliable content shared. We observe two cliques with high similarity: one made up of mainstream platforms (Facebook, Twitter, Reddit) that share a majority of reliable news sources, and one made up of alt-tech ones (Gab, Parler, Voat) sharing a higher fraction of questionable sources. Scored, BitChute, and, to an extent, YouTube remain fairly separated from the rest of the platforms.

### v2

@@ Abstract
The entertainment-driven nature of social media encourages users to engage with like-minded individuals and consume content aligned with their beliefs, limiting exposure to diverse perspectives. Simultaneously, users migrate between platforms, either due to moderation policies like de-platforming or in search of environments better suited to their preferences. These dynamics drive the specialization of the social media ecosystem, shifting from internal echo chambers to “echo platforms”—entire platforms functioning as ideologically homogeneous niches. To systematically analyze this phenomenon in political discussions, we propose a quantitative approach based on three key dimensions: platform centrality, news consumption, and user base composition. We analyze 117 million posts related to the 2020 US Presidential elections from nine social media platforms—Facebook, Reddit, Twitter, YouTube, BitChute, Gab, Parler, Scored, and Voat. Our findings reveal significant differences among platforms in their centrality within the ecosystem, the reliability of circulated news, and the ideological diversity of their users, highlighting a clear divide between mainstream and alt-tech platforms. The latter occupy a peripheral role, feature a higher prevalence of unreliable content, and exhibit greater ideological uniformity. These results highlight the key dimensions shaping the fragmentation and polarization of the social media landscape.
1 Introduction
@@
Social media have become a major source of news and opinions for many, reshaping how people access information and engage in public discussions. This shift coincides with a decline in trust in traditional media [1, 2] and is driven by platforms that prioritize entertainment and user engagement over informational accuracy [3, 4, 5, 6]. In this context, the combination of recommendation algorithms, designed to maximize engagement, and user preferences may encourage individuals to join tightly-knit groups of like-minded peers and consume mostly content that aligns with their pre-existing beliefs, reinforcing shared perspectives and potentially limiting exposure to diverse viewpoints [7, 5, 8]. This behavior has been closely linked to heightened exposure to hate speech [9, 10, 11, 12], political polarization [13, 14, 15], and the formation of ideologically homogeneous clusters of users, called echo chambers [5, 16, 17, 18, 19]. Over time, the echo chamber effect may drive groups towards more extreme positions [20, 21] and influence user behavior across platforms [5, 22, 23] going even beyond digital interactions [24, 25].
To address these challenges, platforms have increasingly implemented moderation policies aimed at managing tight-knit, “problematic” communities [26]. However, moderation policies, such as de-platforming, often push users toward minimally regulated platforms [16, 27, 23, 28, 29]. At the same time, users may spontaneously leave platforms where they find limited alignment with their preferences or feel overly exposed to toxic content, seeking alternatives. For example, many users migrated from Twitter to Mastodon [30] or BlueSky [31, 32] following changes in platform management. When users leave mainstream platforms, they often migrate to less popular digital spaces, commonly referred to as “alt-tech” or “fringe” platforms [33, 34, 23]. Examples include Gab, Parler, BitChute, and Rumble, which have emerged as counterparts to mainstream services like Twitter, Reddit, YouTube, and Facebook. These platforms frequently market themselves as champions of free speech, host alt-right or extremist content, and appeal to users who perceive mainstream platforms as ideologically biased or overly regulated. While their user bases are relatively small compared to mainstream platforms, these alt-tech environments can play a non-negligible role in shaping societal events, serving as important hubs for specific ideological groups, facilitating radicalization, amplifying fringe narratives, contributing to online fragmentation, and creating pathways from online discourse to offline action. For example, Parler was used to coordinate the Capitol Hill assault during the 2020 U.S. elections [35, 36]; similarly, Gab was linked to the Pittsburgh synagogue shooter [37], and Voat became a refuge for QAnon communities banned from Reddit [23].
@@
Both spontaneous and forced migrations foster differentiation and specialization within the social media ecosystem, with platforms increasingly organized around specific user communities and shared content preferences [38, 39, 40, 41]. These shifts represent a departure from the original vision of the world wide web as a globally interconnected network where information flows freely across borders [42, 43]. Instead, entire platforms are now defined by ideologically homogeneous communities and aligned content, creating niches where opposing perspectives are rarely encountered [39]. This scenario—where an entire platform is characterized by a like-minded user base with limited exposure to diverse opinions—lays the foundation for the concept of an “echo platform”, an extension of the echo chamber phenomenon to the level of entire digital ecosystems.
@@
In this paper, we introduce three operational steps to characterize the fragmentation of the social media ecosystem, providing a methodological basis for identifying the role of different platforms. Using a dataset comprising over 117 million URLs collected from nine platforms—four mainstream (Facebook, Twitter, Reddit, YouTube) and five alt-tech fringe platforms (BitChute, Gab, Parler, Scored, Voat)—and nearly six million unique users, during the 2020 US presidential election, we define three axes to characterize platform roles: (i) centrality (central vs. peripheral), (ii) news consumption (reliable vs. questionable content), and (iii) user base composition (uniform vs. diverse). In this context, we adopt the terms mainstream and alt-tech following the terminology commonly used in prior literature [27, 34]: we define alt-tech platforms as those that have been reported to systematically host alt-right or extremist content. However, our goal is to move beyond these binary labels, providing a comprehensive data-driven characterization of the social media ecosystem.
@@
Our analysis reveals a distinct pattern of fragmentation characterized by limited interactions. Alt-tech platforms exhibit significantly higher levels of ideological homogeneity, with user communities sharing content almost exclusively aligned with their dominant narratives. Interestingly, user behavior on Reddit more closely resembles that of alt-tech platforms rather than mainstream ones, featuring a uniformly composed user base that is, however, left-leaning, and playing a relatively peripheral role in cross-platform linking. Furthermore, alt-tech platforms disproportionately amplify questionable content while showing a notable absence of reliable news sources compared to their mainstream counterparts.
@@ Results
The systematic characterization of these three dimensions provides a framework for understanding the role of different platforms in the social media ecosystem. Our analysis is based on a comprehensive dataset of URLs collected from nine social media platforms.
@@
Facebook 300k 10.7M 74k 0.07 0.15 0.16
Reddit 59k 320k 9k 0.10 0.03 0.04
Twitter 5M 103M 114k 0.14 0.16 0.16
YouTube 12k 512k 12k 0.06 0.14 0.19
BitChute 2k 107k 4k 0.18 0.22 0.17
Gab 15k 453k 7k 0.16 0.65 0.06
Parler 73k 1.2M 13k 0.10 0.63 0.06
Scored 41k 868k 11k 0.04 0.78 0.04
Voat 20k 135k 6k 0.13 0.42 0.10
As shown in Table 1, we include four “mainstream” platforms (Facebook, Twitter, Reddit, and YouTube) and five “alt-tech” platforms (BitChute, Gab, Parler, Scored, and Voat). For each platform, we collect only politically relevant posts, especially those related to the 2020 U.S. presidential campaign. Such content was obtained by querying for keywords related to the presidential candidates contained in the posts (Facebook, Twitter, Gab, and Parler) or the video titles or descriptions (YouTube and BitChute). On Reddit, Voat, and Scored posts are often made of only URLs linking to the relevant piece of news, thus a keyword search would not be applicable. As such, we only selected content coming from political communities active during the 2020 U.S. presidential campaign, after a manual inspection of the posts. The time window for the analysis spans year 2020, with different months covered for each platform depending on data availability as detailed in the Materials and Methods section. While we report results based on the full data sets, we provide the analysis using uniformly restricted time frames in Figure S5-S9 of the Supplementary Information (SI), showing that the results remain consistent.
@@
To analyze the roles of different platforms within the social media ecosystem, we model it as a weighted, directed graph. In this network, nodes represent platforms, and an edge from platform $i$ to platform $j$ exists if an account of platform $i$ posted a URL linking directly to platform $j$. The edge weight represents the number of URL links pointing from one platform to another. Therefore, the weighted adjacency matrix of this network (see SI, Fig. S1) can provide insights into how platforms redirect users and share content within the ecosystem. To account for variations in the size of the datasets collected from each platform (and their popularity), we divide the observed weights by the weights expected in a random network which preserves the total number of outgoing and incoming URL links for each platform (see section Materials and Methods for details).
Fig. 1 displays the ratio of observed edge weights to those expected under the null model. Ratios greater than one indicate that a platform links to another more frequently than expected in a random network, while ratios less than one indicate less frequent linking.
@@
Fig. 1 reveals distinct patterns of inter-platform linking: Twitter consistently receives more URLs than expected from all other platforms (third column), positioning it as a central node in the social media ecosystem. Surprisingly, Facebook receives fewer links than expected from other mainstream platforms (first column), except YouTube, likely due to the prevalence of self-promotional links in video descriptions. Alt-tech platforms generally receive fewer links than expected by mainstream ones, and are instead more tightly connected among themselves. Despite this observation, the third highest weight with respect to the null model is observed in the edge from Gab to Reddit, suggesting a potential content-sharing dynamic or migration pathway between these two platforms. In June 2020, indeed, Reddit banned the subreddit r/The_Donald, a forum dedicated to President Trump’s supporters with nearly $800,000$ users, due to repeated violations of the platform’s rules against harassment, hate speech, and content manipulation [44]. Given the substantial presence of Trump supporters on Gab, this could suggest that links from Gab to Reddit were possibly posted as a means of mockery or demotion.
BitChute is linked much more than expected by all fringe platforms and slightly less by mainstream ones (fifth column), while the opposite is true for YouTube (fourth column). Furthermore, we observe a strong relationship between BitChute and Gab, as indicated by the two edges having the highest weights in the network. Finally, Scored does not receive incoming links from six of the eight other platforms, likely due to its limited popularity and low exposure, underscoring its peripheral role in the broader ecosystem.
To better quantify the relative importance of each platform within the ecosystem, we computed their PageRank centrality from the rescaled adjacency matrix shown in Fig. 1. The results presented in Table 1 reveal that BitChute, Gab, and Twitter exhibit the highest scores, reflecting their prominent roles in the network. Voat, Parler, and Reddit display lower centrality scores but still maintain relatively strong connection patterns. Facebook, YouTube, and Scored, on the other hand, rank as the least central platforms in the ecosystem. Therefore, after accounting for popularity, we found that not all mainstream platforms are central, while fringe platforms like Gab and BitChute are not necessarily peripheral.
@@
Next, we investigate whether the type of news and content shared differs significantly across the nine platforms by analyzing the news domains’ political bias. To this aim, we use Media Bias/Fact Check (commonly referred to as MBFC https://mediabiasfactcheck.com), a widely referenced news rating agency that provides political bias and reliability labels for news outlets and other information producers [45, 5, 46]. We extract the domains from the URLs shared across the nine platforms and match them with the political bias and reliability information obtained from MBFC. Domains not classified by MBFC are labeled as “unreported”, though this designation does not necessarily imply an absence of political bias. Following previous research [46], domains categorized as “extreme-left” are excluded from the analysis due to their negligible presence in the dataset (see Materials and Methods for details regarding the labeling of news sources and SI, Table S5, for a table showing the percentage of domains matched with a MBFC rating).
Fig. 2 illustrates users’ news consumption across platforms, showing the proportion of URLs directing to domains with differing political leanings, as categorized by MBFC. A clear pattern emerges: mainstream platforms such as Facebook, Reddit, and Twitter tend to skew toward left-leaning content, while alt-tech platforms—including Gab, Parler, Voat, and Scored—predominantly link to right-leaning content. Voat stands out among alt-tech platforms by directing a notable portion of its traffic to left- or left-center-leaning domains. However, 89.5% of these links were shared across 6 of the 60 distinct subverses analyzed, most of which (e.g., v/AnonAll, v/GreatAwakening) align with alt-right or conspiracy groups (a detailed list of these subverses is available in Table S4 of the SI). This pattern suggests that these links may have been shared in an ironic or derogatory manner. YouTube and BitChute, despite being filtered for political content, heavily link to “unreported” domains, reflecting their role as platforms that host a broad range of less conventional or unvetted content. Additionally, on BitChute, a substantial share of content from hyper-extremist sites is driven by a small group of highly active users. In some cases, these users are only a few dozen or hundreds, yet they are responsible for thousands of links, creating a disproportionate volume of extremist content on the platform. This finding aligns with the concept of “vocal minorities” [47], where a small but highly active subset of users exerts an outsized influence on the platform by driving a significant share of its activity [48].
@@
To further illustrate these patterns, Table 2 reports each platform’s ten most frequently linked domains. While Facebook and Twitter often feature sources with a left-center bias, such as CNN, The New York Times, or The Washington Post, Reddit displays a more homogenous profile as no right-biased sources appear among the most linked domains. In contrast, platforms such as Gab, Parler, Scored, and Voat direct a substantial proportion of their traffic to sources that are both right-biased and have low reliability. Scored, for instance, channels 58% of its external traffic to “Patriots.Win”, an online forum created following the ban of the “r/The_Donald” subreddit in June 2015 [34]. This forum is known for using unreliable sources and a lack of transparency and moderation [49, 50]. Although both YouTube and BitChute frequently link to unreported domains, BitChute’s links often direct users to alternative news websites, including domains associated with the american alt-right.
@@
Facebook Reddit Twitter YouTube BitChute Domain % links Domain % links Domain % links Domain % links Domain % links CNN 2.35 The Hill 6.00 The New York Times 6.90 Fox News* 8.09 Infowars* 3.14 Fox News* 2.30 CNN 5.82 The Washington Post 6.56 MSNBC 1.97 Liberty Classroom* 2.27 The Hill 1.79 The Washington Post 4.73 CNN 5.72 Tv9Hindi 1.86 Rebel News* 1.65 The Washington Post 1.72 The New York Times 3.95 Fox News* 3.22 The Young Turks 1.77 Tuttle Twins 1.65 The New York Times 1.65 Politico 2.79 Politico 2.34 CNBC 1.55 The Gateway Pundit* 1.56 NBC News 1.57 Reuters 2.37 The Hill 2.30 PBS NewsHour 1.47 TimCast 1.53 Yahoo News 1.50 Associated Press 2.18 NBC News 2.27 Fox Business 1.47 Martin Brodel 1776 1.51 Breitbart* 1.46 The Guardian 2.10 Breitbart* 2.23 NBC News 1.10 PeteLive 1.48 Daily Wire 1.22 NBC News 2.07 Raw Story 2.19 France 24 0.91 TurleyTalks* 1.45 MSN 1.04 Business Insider 1.93 The Gateway Pundit* 1.92 CBS News 0.67 Banned* 1.17 Gab Parler Scored Voat Domain % links Domain % links Domain % links Domain % links The Gateway Pundit* 13.71 The Gateway Pundit* 13.74 Patriots* 58.24 The Gateway Pundit* 7.47 Breitbart* 12.07 Fox News* 11.44 Breitbart* 3.39 Breitbart* 5.58 Left Fox News* 4.09 Breitbart* 8.56 MAGA* 2.64 Zero Hedge* 4.91 Left Center Zero Hedge* 2.47 The Epoch Times* 4.94 The Gateway Pundit* 2.33 Fox News* 2.49 Center The Epoch Times* 2.43 Western Journal* 2.44 Fox News* 1.65 DailyMail* 1.92 Right Center InfoWars* 2.29 New York Post 1.74 Zero Hedge* 0.99 The Hill 1.91 Right GNews* 1.97 NewsMax* 1.32 New York Post 0.92 RT News* 1.73 Extreme Right New York Post 1.70 Town Hall* 1.31 Washington Examiner* 0.73 New York Post 1.57 Unreported The Hill 1.31 Just The News* 1.26 DailyMail* 0.70 CNN 1.44 DailyCaller 1.19 CounterGlobalist 1.16 DailyCaller 0.56 Washington Examiner 1.35
Next, we analyze the similarity between the news diets across platforms. To this aim, we identify the top 20 most-linked domains for each platform and compute the weighted cosine similarity between the domain vectors for all pairs. This method quantitatively measures how closely platforms align in their shared content.
The results, illustrated in Fig. 3, reveal two distinct cliques characterized by high similarity scores. The first clique consists of the mainstream platforms Facebook, Twitter, and Reddit, all exhibiting cosine similarity scores above $0.7$. The second clique includes the alt-tech platforms Gab, Voat, and Parler, with similarity scores ranging from $0.73$ to $0.88$. See SI for the full cosine similarity matrix and for a robustness check expanding the threshold of the top 20 most-linked domains (Fig. S2, S3).
@@
The similarity between mainstream and alt-tech platforms generally does not exceed $0.5$, with Facebook and Twitter showing slightly higher similarity to specific fringe platforms. This last observation may indicate that news consumption withing certain echo chambers on Facebook and Twitter partially overlaps with those on fringe platforms. Reddit appears to be the mainstream platform most dissimilar to alt-tech platforms, exhibiting the lowest similarity in news consumption. This aligns with the tendency of Reddit users to predominantly share left-leaning content, while extreme-right sources are absent and right-center or right-leaning content is relatively uncommon. YouTube shows moderate similarity to other mainstream platforms, particularly Twitter, but also with some alt-tech ones, such as Parler. Finally, BitChute and Scored are dissimilar to all other platforms. This difference may be attributed to the extreme nature of the content frequently shared on BitChute and the limited popularity and reach of Scored, which results in unique news diets that diverge significantly both from mainstream and other alt-tech platforms.
@@
Fig. 3 also highlights the fraction of questionable domains (as defined by MBFC) shared by each platform. Mainstream platforms share a relatively small fraction of unreliable sources, while alt-tech platforms display significantly higher proportions of such content, reinforcing the distinct content dynamics within the social media ecosystem. These results are reported in more detail in Table 1, showing the fraction of questionable sources shared by users on each platform. Alt-tech platforms Gab, Parler, and especially Scored exhibit the highest proportions. Voat shows a lower fraction, and Reddit stands out for its smallest fraction of questionable sources shared ($0.03$). These findings underscore the distinction between alt-tech platforms, where questionable sources appear to circulate more freely, and mainstream platforms, where such content constitutes only a small proportion of shared material. BitChute shows a small value of $0.22$, differing from the general trend observed among alt-tech platforms, but aligning with the platform’s propensity to host a significant volume of unreported domains.
@@
Figure 3: Platform similarity. Cosine similarity network based on the platforms’ 20 most linked domains. The size of the different nodes is proportional to the volume of links shared in the platform, while the colors of the pies indicate the fraction of questionable or reliable content shared. We observe two cliques with high similarity: one made up of mainstream platforms (Facebook, Twitter, Reddit) that share a majority of reliable news sources, and one made up of alt-tech ones (Gab, Parler, Voat) sharing a higher fraction of questionable sources. Scored, BitChute, and, to an extent, YouTube remain fairly separated from the rest of the platforms.